I am not sure this problem is a Polynomial,any help you can offer would be great. Sorry the -2 is under the 1st set of parenthsis & -4 is under the 2nd set. Thanks
4+ (-3x+4)=(-2x+6)+24
------------
-2 -4



Answer by nyc_function(2741)

Do you mean the following?
[4 + (-3x + 4)]/-2 = [(-2x+6)+ 24]/-4
Is this what you are trying to simplify?
Simplify numerator as step one.
(-3x + 8)/-2 = (-2x + 30)/-4
Multiply both sides of the equation by the LCD -4. This multiplication process will remove the fractions from both sides of the fractional equation.
(-2)(-3x + 8) = -2x + 30
6x - 16 = -2x + 30
Can you solve for x now?